# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v4

@@ 2. Description of the set of canonical forms
Proof. Fix a root $\gamma\in D$. Consider the orbit of the infinitesimal action $\mathfrak{n}.\lambda$. Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, where $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. Here $\mu$ is given, while the values $x_{\alpha}$ are unknown.
@@
$\lambda([e_{\delta},e_{\beta}])x_{\delta}=-\mu(e_{\beta})-\sum\limits_{\alpha\neq\delta}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
@@
[…]
Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ will also be satisfied for it. Therefore, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none do. Hence, it suffices to check the conditions for a single form with a given support, for example, for the form with all coordinates equal to 11 at the basis vectors $e_{\alpha}^{*}$ with $\alpha\in D$.
Proposition 2.4. (Necessary condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$, $\gamma\in D$. Suppose that there exist positive roots $\delta_{1},\delta_{2},\dots\delta_{k},\epsilon_{1},\epsilon_{2},\dots\epsilon_{k}$ (not necessarily distinct) satisfying the conditions:
@@
    $\lambda([e_{\delta_{i}},e_{\beta}])\neq 0$ and $\beta\notin\zeta_{D,\gamma}\iff\beta=\epsilon_{i-1}$ or $\beta=\epsilon_{i}$ for all $i=\overline{1,k-1}$;
@@
    $\lambda([e_{\alpha},e_{\epsilon_{i}}])\neq 0$ and $\alpha\succeq\gamma\iff\alpha=\delta_{i}$ or $\alpha=\delta_{i+1}$ for all $i=\overline{1,k}$;
@@
    $\lambda([e_{\delta_{k}},e_{\epsilon_{k}}])\neq 0$ and $\epsilon_{k}\notin\zeta_{D,\gamma}$.
@@
Proof. It is enough to show that $(x.\lambda)(e_{\gamma})$ cannot be expressed as a linear combination of $\{(x.\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. In other words, consider the orbit of the tangent action $\mathfrak{n}.\lambda$, and let $\mu\in\mathfrak{n}.\lambda$. It is enough to check that if all values $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary. We will construct $\mu^{\prime}\in\mathfrak{n}.\lambda$ such that $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is an arbitrary given value. Let $\mu^{\prime}=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, to be the same as for $\mu$. We will determine the values $x_{\epsilon_{i}}$ by induction on $i=\overline{1,k-1}$.
@@
Base case. Let $i=1$, $A=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$. Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$, then either $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$. Hence, in the right-hand side, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ except possibly $x_{\epsilon_{1}}$, are already defined. Since $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined as well. After this, the only expression $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that may still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$. If $k=1$, then nothing else depends on $x_{\epsilon_{1}}$.
@@
Inductive step. Suppose the value $x_{\epsilon_{i-1}}$ has been found. Consider $\mu(e_{\delta_{i}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$. On the right-hand side, only $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$, can appear with non-zero coefficients. The value $x_{\epsilon_{i-1}}$ was determined in the previous step. Thus, on the right-hand side, only $x_{\epsilon_{i}}$ may be undefined, and it definitely appears. From this equation, the value $x_{\epsilon_{1}}$ is uniquely determined. The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that may still depend on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
@@
Case $i=k$. Consider $\mu(e_{\delta_{k}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$. The right-hand side definitely contains $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$. The value $x_{\varepsilon_{k-1}}$ was determined in the previous step (or does not appear if $k=1$). All other $x_{\alpha}$ are also defined. Therefore, the value $x_{\epsilon_{k}}$ is uniquely determined. Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depends on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
@@
Thus, we have computed all coefficients $x_{\alpha}$ and constructed the required $\mu^{\prime}$. $\square$
@@
Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the specific coordinates of $\lambda$, but only on its support. Therefore, it is sufficient to check these conditions for just one form with a given support; for example, we may assume that all its nonzero coordinates are equal to 11.
@@
The conditions of Propositions Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked algorithmically. Below we present algorithms for checking the necessary and sufficient conditions for a support to belong to the set SS."
@@
The enumeration proceeds according to the following rules. First, the support $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$} is checked against the necessary (respectively, sufficient) condition. Next, suppose we have checked a support $D$ against the necessary (respectively, sufficient) condition. Let $\delta$ be the minimal root in $D$. First, let us consider the case if $\delta\neq\alpha_{4}$. If $D$ passes the check, we add to $D$ the positive root immediately preceding $\delta$ (with respect to the order $\succ$). If $D$ fails the check, we remove $\delta$ from $D$ and add the positive root immediately preceding $\delta$. Then we check the new support against the necessary (respectively, sufficient) condition. Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if it exists) and add the root closest to $\beta$ from below. If $D=\{\alpha_{4}\}$, then the enumeration ends here. Finally, we include the empty set in the list of supports, since it trivially satisfies the necessary (respectively, sufficient) condition. In this way, we enumerate all supports that could potentially satisfy the necessary (respectively, sufficient) condition, while avoiding checks for supports that certainly do not satisfy these conditions.
@@
Checking the sufficient condition for the support $D$ is straightforward. We iterate over the row indexed by $\gamma$ and verify the condition from Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
@@
Linear forms with support 5 do not satisfy the necessary condition, but checking this requires three steps, whereas our algorithm checks only the first two steps (see above). It is clear that linear forms with supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1); that is, such forms do not belong to $S$. This follows immediately from Table 1, similar to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. In the remaining cases, it is necessary to write out the coordinates of the infinitesimal action explicitly. It turns out that the linear forms with supports numbered 25 and 33 lie in $S$ without any restrictions on their coordinates. Linear forms with supports numbered 30, 31, and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$. Thus, we obtain 883 supports whose corresponding forms belong to $S$, three of which require an additional condition on the coordinates. This completes the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
@@
Example 2.5. Consider a form $\lambda$ with support 33. We will show that $\lambda\in S$. Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}^{*}$. Let us write down explicitly the coordinates of interest. We denote the structure constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. Then
$\begin{split}\mu_{\alpha_{1}+\alpha_{2}}&=\mu_{\alpha_{2}+2\alpha_{3}}=\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=0;\\ \mu_{\alpha_{1}}&=-\lambda([x,e_{\alpha_{1}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{1}}])=-C_{\alpha_{2},\alpha_{1}}x_{\alpha_{2}}\lambda(e_{\alpha_{1}+\alpha_{2}})=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ &=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\lambda_{\alpha_{1}+\alpha_{2}}}\mu_{\alpha_{1}};\\ \mu_{\alpha_{3}}&=-\lambda([x,e_{\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{3}+\alpha_{4}}-2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{4}}&=-\lambda([x,e_{\alpha_{4}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])\\ &=-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\\ &=\dfrac{\lambda_{\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{2}+\alpha_{3}}-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{3}}.\end{split}$
@@
We see that if $\beta\succ\alpha_{2}$ and the variable $x_{\alpha_{3}+\alpha_{4}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+\alpha_{3}$. Similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+2\alpha_{3}$. Hence, the coordinate $\mu_{\alpha_{2}}$ must be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$. It is easy to verify that this holds if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
@@
By an argument analogous to the one above, we find that the coordinate $\mu_{\alpha_{4}}$ must be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$. It turns out that this condition is also equivalent to $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
3. Classification of orbits
In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, thereby completing the classification of coadjoint orbits for type $F_{4}$. For the next stage, we will need the concept of a singular root.
Definition 3.1. Let $\alpha,\gamma\in\Phi^{+}$. The root $\alpha$ is called singular to $\gamma$ if $\gamma-\alpha\in\Phi^{+}$. The set of all roots singular to $\gamma$ is denoted by $S_{\gamma}$.

### v5

@@ 2 Description of the set of canonical forms
Fix a root $\gamma\in D$. Consider the orbit of the infinitesimal action $\mathfrak{n}\cdot\lambda$. Let $\mu\in\mathfrak{n}\cdot\lambda$, $\mu=x\cdot\lambda$, where $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. Here $\mu$ is given, while the values $x_{\alpha}$ are unknown.
@@
$\lambda([e_{\delta},e_{\beta}])x_{\delta}=-\mu(e_{\beta})-\sum_{\alpha\neq\delta}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
@@
Now consider $\mu(e_{\gamma})=-\sum_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\gamma}])x_{\alpha}$. By the hypothesis of the proposition, if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Consequently, the same holds for $\mu(e_{\gamma})$, which is equivalent to condition (1) for the root $\gamma$. The result follows.
Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition 2.3 will also be satisfied for it. Therefore, for a given support $D$, either all forms with this support satisfy the conditions of Proposition 2.3, or none do. Hence, it suffices to check the conditions for a single form with a given support, for example, for the form with all coordinates equal to 11 at the basis vectors $e_{\alpha}^{*}$ with $\alpha\in D$.
Proposition 2.5.
(Necessary condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$, $\gamma\in D$. Suppose that there exist positive roots $\delta_{1},\delta_{2},\dots\delta_{k},\varepsilon_{1},\varepsilon_{2},\dots\varepsilon_{k}$ (not necessarily distinct) satisfying the conditions:
@@
    $\lambda([e_{\delta_{i}},e_{\beta}])\neq 0$ and $\beta\notin\zeta_{D,\gamma}\iff\beta=\varepsilon_{i-1}$ or $\beta=\varepsilon_{i}$ for all $i=\overline{1,k-1}$;
@@
    $\lambda([e_{\alpha},e_{\varepsilon_{i}}])\neq 0$ and $\alpha\succeq\gamma\iff\alpha=\delta_{i}$ or $\alpha=\delta_{i+1}$ for all $i=\overline{1,k}$;
@@
    $\lambda([e_{\delta_{k}},e_{\varepsilon_{k}}])\neq 0$ and $\varepsilon_{k}\notin\zeta_{D,\gamma}$.
@@
It is enough to show that $(x\cdot\lambda)(e_{\gamma})$ cannot be expressed as a linear combination of $\{(x\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. In other words, consider the orbit of the tangent action $\mathfrak{n}\cdot\lambda$, and let $\mu\in\mathfrak{n}\cdot\lambda$. It is enough to check that if all values $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary. We will construct $\mu^{\prime}\in\mathfrak{n}\cdot\lambda$ such that $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is an arbitrary given value. Let $\mu^{\prime}=x\cdot\lambda$, $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. It follows from the proof of Proposition 2.3 that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. We define all other values $x_{\alpha}$, except $x_{\varepsilon_{i}}$, to be the same as for $\mu$. We will determine the values $x_{\varepsilon_{i}}$ by induction on $i=\overline{1,k-1}$.
@@
Base case. Let $i=1$, $A=\sum_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$. Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$, then either $\alpha=\varepsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$. Hence, in the right-hand side, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ except possibly $x_{\varepsilon_{1}}$, are already defined. Since $\lambda([e_{\delta_{1}},e_{\varepsilon_{1}}])\neq 0$, the value of $x_{\varepsilon_{1}}$ is uniquely determined as well. After this, the only expression $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that may still depend on $x_{\varepsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$. If $k=1$, then nothing else depends on $x_{\varepsilon_{1}}$.
@@
Inductive step. Suppose $x_{\varepsilon_{i-1}}$ has been found. Consider $\mu(e_{\delta_{i}})=\sum_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$. On the right-hand side, only $x_{\varepsilon_{i-1}},x_{\varepsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$, can appear with non-zero coefficients. The value $x_{\varepsilon_{i-1}}$ was determined in the previous step. Thus, on the right-hand side, only $x_{\varepsilon_{i}}$ may be undefined, and it definitely appears. From this equation, the value $x_{\varepsilon_{1}}$ is uniquely determined. The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that may still depend on $x_{\varepsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
@@
Case $i=k$. Consider $\mu(e_{\delta_{k}})=\sum_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$. The right-hand side definitely contains $x_{\varepsilon_{k}}$ and possibly $x_{\varepsilon_{k-1}}$. The value $x_{\varepsilon_{k-1}}$ was determined in the previous step (or does not appear if $k=1$). All other $x_{\alpha}$ are also defined. Therefore, the value $x_{\varepsilon_{k}}$ is uniquely determined. Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depends on $x_{\varepsilon_{k}}$, and its value no longer affects anything else.
@@
Thus, we have computed all coefficients $x_{\alpha}$ and constructed the required $\mu^{\prime}$.
@@
Also note that the conditions of Proposition 2.5 do not depend on the specific coordinates of $\lambda$, but only on its support. Therefore, it is sufficient to check these conditions for just one form with a given support; for example, we may assume that all its nonzero coordinates are equal to 11.
@@
The conditions of Propositions 2.3 and 2.5 are formulated so that they can be checked algorithmically. Below we present algorithms for checking the necessary and sufficient conditions for a support to belong to the set $S$.
@@
The enumeration proceeds according to the following rules. First, the support of the element $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$} is checked against the necessary (respectively, sufficient) condition. Next, suppose we have checked a support $D$ against the necessary (respectively, sufficient) condition. Let $\delta$ be the minimal root in $D$. First, let us consider the case if $\delta\neq\alpha_{4}$. If $D$ passes the check, we add to $D$ the positive root immediately preceding $\delta$ (with respect to the order $\succ$). If $D$ fails the check, we remove $\delta$ from $D$ and add the positive root immediately preceding $\delta$. Then we check the new support against the necessary (respectively, sufficient) condition. Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if it exists) and add the root closest to $\beta$ from below. If $D=\{\alpha_{4}\}$, then the enumeration ends here. Finally, we include the empty set in the list of supports, since it trivially satisfies the necessary (respectively, sufficient) condition. In this way, we enumerate all supports that could potentially satisfy the necessary (respectively, sufficient) condition, while avoiding checks for supports that certainly do not satisfy these conditions.
@@
Checking the sufficient condition for the support $D$ is straightforward. We iterate over the row indexed by $\gamma$ and verify the condition from Proposition 2.3.
@@
Linear forms with support 5 do not satisfy the necessary condition, but checking this requires three steps, whereas our algorithm checks only the first two steps (see above). It is clear that linear forms with supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1); that is, such forms do not belong to $S$. This follows immediately from Table 1, similar to Example $2.$ In the remaining cases, it is necessary to write out the coordinates of the infinitesimal action explicitly. It turns out that the linear forms with supports numbered 25 and 33 lie in $S$ without any restrictions on their coordinates. Linear forms with supports numbered 30, 31, and 32 lie in $S$ if and only if
Thus, we obtain 883 supports whose corresponding forms belong to $S$, three of which require an additional condition on the coordinates. This completes the proof of Proposition 2.1.
Example 2.7.
@@
Consider a form $\lambda$ with support 33. We will show that $\lambda\in S$. Let $\mu\in\mathfrak{n}\cdot\lambda$, $\mu=x\cdot\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}^{*}$. Let us write down explicitly the coordinates of interest. We denote the structure constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [10]. Then
$\begin{split}\mu_{\alpha_{1}+\alpha_{2}}&=\mu_{\alpha_{2}+2\alpha_{3}}=\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=0;\\ \mu_{\alpha_{1}}&=-\lambda([x,e_{\alpha_{1}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{1}}])=-C_{\alpha_{2},\alpha_{1}}x_{\alpha_{2}}\lambda(e_{\alpha_{1}+\alpha_{2}})=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=\\ &=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\\ &=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ &=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\lambda_{\alpha_{1}+\alpha_{2}}}\mu_{\alpha_{1}};\\ \mu_{\alpha_{3}}&=-\lambda([x,e_{\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\\ &-x_{\alpha_{4}}\lambda_{\alpha_{3}+\alpha_{4}}-2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{4}}&=-\lambda([x,e_{\alpha_{4}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])\\ &=-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\\ &=x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\dfrac{\lambda_{\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{2}+\alpha_{3}}-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{3}}.\end{split}$
@@
We see that if $\beta\succ\alpha_{2}$ and the variable $x_{\alpha_{3}+\alpha_{4}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+\alpha_{3}$. Similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+2\alpha_{3}$. Hence, the coordinate $\mu_{\alpha_{2}}$ must be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$. It is easy to verify that this holds if and only if
@@
By an argument analogous to the one above, we find that the coordinate $\mu_{\alpha_{4}}$ must be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$. It turns out that this condition is also equivalent to
3 Classification of orbits
In this section, we will prove Theorem 1.1, thereby completing the classification of coadjoint orbits for type $F_{4}$. For the next stage, we will need the concept of a singular root.
Definition 3.1.
Let $\alpha,\gamma\in\Phi^{+}$. The root $\alpha$ is called singular to $\gamma$ if $\gamma-\alpha\in\Phi^{+}$. The set of all roots singular to $\gamma$ is denoted by $S_{\gamma}$.